OpenStudy (anonymous):

so is there a certain formula i use to set this up A brownie recipe requires one and one third cups of sugar to one cup of chocolate chips. If three and one third cups of sugar is used, what quantity of chocolate chips will be needed, according to the recipe?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

easy, just setup a proportion

ganeshie8 (ganeshie8):

proportion involves two ratios :- first ratio : one and one third cups of sugar to one cup of chocolate chips. \(\large \frac{1\frac{1}{3}}{1}\)

ganeshie8 (ganeshie8):

second ratio : three and one third cups of sugar is used \(\large \frac{3\frac{1}{3}}{x} \)

ganeshie8 (ganeshie8):

set them equal and solve x

OpenStudy (anonymous):

i did wrong ? lol i meant to put this one A pancake recipe requires two and one half cups of milk to one cup of flour. If three and one third cups of milk is used, what quantity of flour will be needed, according to the recipe?

ganeshie8 (ganeshie8):

its same, just setup a proportion. dont get distracted by the mixed numbers... this is really simple

ganeshie8 (ganeshie8):

two and one half cups of milk to one cup of flour \(\large \frac{2\frac{1}{2}}{1} \)

ganeshie8 (ganeshie8):

three and one third cups of milk is used \(\large \frac{3\frac{1}{3}}{x} \)

ganeshie8 (ganeshie8):

set both the ratios equal and solve x

ganeshie8 (ganeshie8):

\(\large \frac{2\frac{1}{2}}{1} = \frac{3\frac{1}{3}}{x} \)

ganeshie8 (ganeshie8):

\(\large x = \frac{3\frac{1}{3}}{2\frac{1}{2}} \)

OpenStudy (anonymous):

so 1=1/x ?

ganeshie8 (ganeshie8):

\(\large x = \frac{\frac{10}{3}}{\frac{5}{2}} \)

ganeshie8 (ganeshie8):

\(\large x = \frac{10}{3} \times \frac{2}{5} \)

ganeshie8 (ganeshie8):

\(\large x = \frac{4}{3} \)

ganeshie8 (ganeshie8):

\(\large x = 1\frac{1}{3} \)

ganeshie8 (ganeshie8):

so you will need one and one thirds of chocolate chips

ganeshie8 (ganeshie8):

go slow, and see if it makes sense :)

OpenStudy (anonymous):

im trying a new problem i think i understand from what u showed me (: if i need help ill post it(:

ganeshie8 (ganeshie8):

sounds good :)